# Mixed Diffusion for 3D Indoor Scene Synthesis

### v1

@@ 4.2 Denoising Network
Figure 1: Denoising network. The time-variant decoder backbone takes object features as input, conditioned by the floor plan feature after concatenation with positional embeddings of object indices. We utilize the time-variant transformer decoder block from VQ-Diffusion [11].
We use a transformer-based denoising network as shown in Fig. 1, which takes in latent variables $(z_{t},{\bm{x}}_{t})$ at step $t$ and predicts the categorical distribution $p_{\theta}(z_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}})$, together with the Gaussian mean of $p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t},z_{t},{\bm{y}})$ assuming a fixed covariance. At test time, we can simply sample from these distributions to retrieve $(z_{t-1},{\bm{x}}_{t-1})$.
Feature encoder. We encode the object features by passing geometric attributes through an MLP and combine them with trainable semantic embeddings. For each scene, we sample 256 points from a floor plan image along the boundary and compute outward normal vectors. We use the PointNet [31] floor plan feature extractor from LEGO-Net [46] as it is more lightweight and better captures floor boundary than image-based feature extractors. The floor plan feature is then concatenated with the learned position embedding, which is commonly used in transformer architectures, to form the conditioning vector.
Transformer decoder. The backbone of our denoising network is a time-variant transformer decoder adapted from VQ-Diffusion [11]. We show in experimentally that the multi-head attention mechanism in the transformer decoder better captures boundary constraints than the conditional 1-D U-Net [34] architecture proposed in DiffuScene [41]. We show our denoising network in Fig. 1. Within each transformer block, the time step is injected through the Adaptive Layer Normalization (AdaLN) [3] operator. The conditioning inputs are sent through a multi-head cross-attention layer.
Feature extractor. We feed the output of the transformer decoder to two MLPs to generate categorical distribution for the semantic labels and Gaussian means for the geometric features.
@@ 4.3 Partial Condition on Objects
3D scene synthesis is sometimes constrained by known object information, such as scene completion given existing objects. DiffuScene [41] considers these as additional condition inputs to the denoising network, which requires at least a task-specific feature encoder for the condition inputs and hence re-training of the entire model. We instead propose to address this flexibility issue through a corruption-and-masking strategy which does not require any additional training. Since the core idea of diffusion models is to recover real data by learning the inverse of a known corruption process, we can force the constrained sections of the latent variables to follow exactly the inverse of the controlled corruption steps in the denoising process. For example, for scene completion: given $M$ objects, we first compute a sequence of latent variables $\{(z_{t}^{i},{\bm{x}}_{t}^{i})\}_{1\leq i\leq M}$ through the corruption process. This allows us to mask out the first $M$ objects using pre-computed latent variables at time $t$ in the reverse process. These latent variables eventually converge to the known input $\{(z_{0}^{i},{\bm{x}}_{0}^{i})\}_{1\leq i\leq M}$ without additional training. This strategy can generalize to other applications, such as furniture arrangement and label conditioned scene synthesis, by targeting different sections of latent object variables.
@@ 5 Experimental Evaluation
In this section, we present evaluation results comparing MiDiffusion against existing methods and provide ablation studies. We also show that, without re-training, our models can be applied to other tasks such as scene completion and furniture arrangement. Additional results and implementation details can be found in the appendix.
@@
Dataset. We use the common benchmarking dataset, 3D-FRONT [9], for evaluation. This is a synthetic dataset of 18,797 rooms furnished by 7,302 textured objects from 3D-FUTURE [10]. We evaluate on three room types: 4041 bedrooms, 900 dining rooms, and 813 living rooms. We use the same data processing and training/testing data split as [28]. We use rotation augmentations in $90^{\circ}$ increments during training as in  [41] to be consistent with the baselines.
@@
Baselines. We compare our proposed approaches against two state-of-the-art 3D scene synthesis methods: (1) ATISS [28], an autoregressive model with a transformer encoder backbone that generates sequential predictions of objects; (2) DiffuScene [41], a diffusion model with a conditional 1D U-Net [34] denoising backbone based on DDPM. We train baseline models with their open-source code and default hyper parameters. We compare MiDiffusion with DiffuScene in its original application, i.e., unconditioned scene synthesis, in Appendix E.2.
@@
Implementation. We train all our models with the Adam optimizer [21] at $l_{r}=2e^{-4}$ with a $0.5$ decay rate every $10k$ for $50k$ epochs on the bedroom dataset, and every $15k$ epochs for $10k$ epochs on the dining room and living room datasets. We set $\lambda=0.05$ for the auxiliary loss term $L_{aux}^{D3PM}$. More details about implementation and network parameters are included in Appendix C.
@@
Evaluation Metrics. We randomly select 1000 floor plans from the test set as input and evaluate models using metrics found in prior works [44, 28, 41]: KL divergence of object label distributions (KL×$\times 0.01$), Fréchet inception distance (FID), Kernel inception distance (KID×$\times 0.001$), and classification accuracy (CA%) using the implementation by ATISS. While the KL divergence evaluates the predicted object label distributions, the other metrics are used to compare the layout images with the ground-truth test data. We render each scene to a 256 x 256 top-down orthographic projection image using CAD models from 3D-FUTURE. In line with DiffuScene, we color each object according to to its semantic class. We render bedrooms to 6.1m squares, and the living and dining rooms to 12.2m squares. Appendix D includes example layout images. We report CA % over 10 runs with random sampling of predicted layouts for training and testing. Note that realistic layouts should have a low FID/KID, a low KL and a CAthat is close to 50%. We also compute additional metrics, averaged over all scenes to further evaluate object placement, including the number of generated objects (Obj), percentage of out-of-boundary objects (OOB %), and bounding box Intersection over Union (IoU %). We dilate the room boundary by 0.1m for OOB % computation to account for slight under estimation of floor boundary in the raw 3D-FRONT data.
@@ 5.1 Floor Plan Conditioned 3D Scene Synthesis
We first compare MiDiffusion against baselines in floor-conditioned 3D scene synthesis experiments without existing objects. Fig. 2 include qualitative results rendered in PyVista. The object mesh is retrieved from the 3D-FUTURE [10] dataset by finding the closest match in size under the predicted semantic category. We replace original textures in the CAD models by label specific color for clarity. We include examples of orthographic projection images, which we use for quantitative evaluations, in the appendix. In general, our approach is able to generate realistic object arrangement whilst respecting the boundary constraints. ATISS generates slightly less precise object placement compared to the two diffusion based approaches, e.g., the table-chairs furniture set in the last example of Fig. 2 is not exactly symmetric. Both DiffuScene and ATISS show higher tendency to place objects outside the floor boundary than ours. This problem is more prominent for DiffuScene, most likely due to the conditional U-Net architecture in their denoising network, which does not learn boundary constraints equally well as transformer architectures.
We show quantitative results in Tables 1 and 2. Overall, our approach achieves substantially better performance in generating realistic layouts, especially in the dining room and living room datasets. There are large margins between our models and the baseline models for FID, KID×$\times 0.001$ and CA %, comparing images of synthesized layouts. As for object label distributions, there seems to be no large differences between the individual methods with respect to KL×$\times 0.01$. In fact, all methods have rather small KL values and the difference could be due to random sampling. In Table 2, there is a trade-off between keeping objects inside the boundary (low OOB%) and maintaining minimal overlap (low IoU%). Although DiffuScene achieves the lowest IoU%, their models tend to place more objects outside the boundary than the others, which explains their sub-optimal performance in Table 1. Overall, our approach is still the best in terms of object placement that are compatible with other predicted objects and the floor boundary constraints.
@@
ATISS 65.76 1.21 54.44 $\pm$ 3.69 0.94 41.06 10.54 64.06 $\pm$ 5.63 1.89 38.03 9.20 63.98 $\pm$ 4.60 1.55
DiffuScene 66.45 1.05 58.70 $\pm$ 4.93 3.41 45.88 9.30 65.34 $\pm$ 6.55 1.55 47.91 11.75 73.96 $\pm$ 4.93 1.95
MiDiffusion (Ours) 63.71 0.24 53.65 $\pm$ 2.46 1.05 30.63 2.74 52.10 $\pm$ 1.72 1.90 28.99 1.60 54.69 $\pm$ 1.54 1.54
@@
ATISS 5.52 16.26 1.00 11.43 17.43 1.86 13.00 17.39 1.49
DiffuScene 5.05 8.06 0.41 10.90 28.16 0.56 11.79 33.75 0.43
MiDiffusion (Ours) 5.22 3.91 0.61 10.92 5.77 0.91 12.28 8.71 0.69
@@
Diversity. We study the diversity of the predicted layouts by measuring the average standard deviation (std) of object centroid positions and bounding box sizes. Since vertical positions are highly correlated with semantic labels, which are already evaluated through KL-divergence, the std over positions is averaged over the two planar axes. The results are in Table 3, where the “IB” suffix means results are computed over “in-boundary” objects only to remove bad object placement. However, these numbers are not directly comparable due to large differences in OOBacross approaches. Nevertheless, our approach generates the most diverse object placements that are within the floorplan boundary (i.e., Position-IB). We are slightly less diverse in sizes (i.e., Size-IB) than DiffuScene, but our approach approach shows the least drop in size variety after removing out-of-boundary objects.
@@
ATISS 1.074 1.035 0.709 0.663 1.586 1.504 0.434 0.385 1.730 1.663 0.442 0.409
DiffuScene 1.073 1.059 0.718 0.691 1.596 1.427 0.454 0.423 1.752 1.575 0.482 0.439
MiDiffusion (Ours) 1.073 1.067 0.698 0.680 1.568 1.558 0.414 0.399 1.670 1.667 0.441 0.427
@@
Ablation study. In Table 4, we study the effect of our mixed-diffusion formulation and PointNet floor plan feature extractor.We provide results of modified versions of MiDiffusion: (1) with DDPM and ResNet-18 [12] as floor plan feature extractor as in ATISS and DiffuScene; (2) with DDPM and PointNet. The last row shows our method MiDiffusion (i.e., using our mixed diffusion formulation together with PointNet). The first variant, DDPM+ResNet is similar to DiffuScene, except we use a transformer based denoising network rather than its U-Net counterpart. Note that the respective results already exceed DiffuScene in Table 1, suggesting that the transformer architecture is better at learning boundary constraints. However, there is a large deterioration in KL divergence. Next, while replacing the ResNet floor feature extractor by a more lightweight PointNet extractor brings a significant performance boost in the dining and living room datasets, which contain less training scenes but are of larger room sizes, the performance decreases on the bedroom dataset. It is likely that PointNet works better under complex and limited training data, but suffers from sampling noise during conversion of the floor plan image to points and normals in the opposite situation. Overall, we achieve the best results with our complete formulation in which the mixed diffusion approach avoids lifting the discrete inputs to a continuous space as required by the DDPM formulation. We include additional geometric evaluation results in Appendix E.1
Method Bedroom Dining room Living room
Mixed+PointNet 63.71 0.24 53.65 $\pm$ 2.46 1.05 30.63 2.74 52.10 $\pm$ 1.72 1.90 28.99 1.60 54.69 $\pm$ 1.54 1.54
Table 4: Ablation study for floor-conditioned 3D scene synthesis.
Recall from Sec. 4.3 that we can re-use the same MiDiffusion models for problems involving partial object constraints through a corruption-and-masking strategy. This approach forces the relevant object attributes to follow the reverse of a pre-computed corruption process in the denoising step, without the need of task-specific retraining.
5.2.1 Scene completion
Figure 3: Bedroom completion given a bed.
Figure 4: Living room completion given sofa and corner tables.
We first compare our approach with baselines for scene completion (i.e., scene synthesis conditioned on existing objects). We train the DiffuScene-SC variant, which includes an additional module to learn existing object features as a conditional input, using their suggested hyper-parameters and providing 3 existing objects during training. The number of existing objects can be arbitrary at test time. We use the same models for ATISS and MiDiffusion. We use the proposed corruption-and-masking strategy over MiDiffusion models trained in Section 5.1. Fig. 4 shows a simple example of completing a bedroom design that originally only consists of a single bed. Fig. 4 shows another more complex example for completion of a living space. Note that MiDiffusion is able to generate a natural and symmetric set of furniture layout.
We perform quantitative evaluations in Table 5. This is an easier problem for ATISS and DiffuScene than the previous setup. In particular, ATISS is an autoregressive model and therefore existing objects allow the models to skip the first few iterations reducing the chance for mistakes. DiffuScene-SC is, on the other hand, specially designed and trained for this task. Based results in Table 1, these approaches indeed benefit from using existing objects as anchors. Without any retraining, our approach still outperforms these approaches, when given only 1 existing object. When given 3 objects, our models are still better for the dining room and living room datasets and, however, are outperformed by DiffuScene-SC on the bedroom dataset. Note that the bedroom data has only an average of only 5.22 objects. Therefore, specially trained networks, such as DiffuScene-SC, might be better suited given the very high percentage of pre-existing objects.
@@ 5.2.1 Scene completion
1 obj ATISS 60.59 0.19 1.04 13.52 36.63 7.77 1.79 15.30 34.34 7.15 1.46 16.19
DiffuScene-SC 83.90 19.17 18.34 3.11 34.12 4.82 1.59 12.13 36.32 6.64 1.30 21.74
MiDiffusion (ours) 60.08 -0.03 0.91 3.32 28.56 1.88 1.58 5.29 28.21 1.78 1.47 8.04
3 obj ATISS 51.83 -0.57 0.66 8.89 33.53 5.39 1.60 12.26 31.78 5.39 1.43 13.03
DiffuScene-SC 42.77 -1.08 0.80 5.13 30.58 3.67 0.84 10.04 32.07 4.58 0.57 16.13
MiDiffusion (ours) 52.19 -0.32 1.00 4.05 27.60 1.85 1.72 3.97 27.46 1.91 0.95 7.22
@@
5.2.2 Furniture Arrangement
Figure 5: Different table-chair arrangements.
Figure 6: Different living room furniture placement directions.
MiDiffusion can also be used for furniture arrangement, which refers to a scene synthesis problem conditioned on object labels and bounding box sizes without an initial configuration. We use our proposed corruption-and-masking strategy over these attributes to test our models in this setting, and compare the results against DiffuScene’s furniture arrangement variant (DiffuScene-FA). ATISS models are not well suited for this task, since replacing the predicted object label by constraints breaks the stopping condition in their autoregressive pipeline. Fig. 6 shows a simple example of a dining room with two different chair arrangements, as generated by MiDiffusion. Fig. 6 instead illustrates two opposite placements of the living room section given the same inputs. We show quantitative results in Table 6. Without re-training, our models outperform DiffuScene-FA in all metrics.

### v2

@@ 4.2 Denoising Network
We use a transformer-based denoising network as shown in Fig. 1, which takes in latent variables $(z_{t},{\bm{x}}_{t})$ in step $t$ and predicts the categorical distribution $p_{\theta}(z_{t-1}|z_{t},{\bm{x}}_{t},{\bm{y}})$, together with the Gaussian mean of $p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t},z_{t},{\bm{y}})$ assuming a fixed covariance. At test time, we can simply sample from these distributions to retrieve $(z_{t-1},{\bm{x}}_{t-1})$.
Feature encoder. We encode the object features by passing geometric attributes through an MLP and combine them with trainable semantic embeddings. For each scene, we sample 256 points from a floor plan image along the boundary and compute outward normal vectors. We use the PointNet [35] floor plan feature extractor from LEGO-Net [51] as it is more lightweight and better captures floor boundary than image-based feature extractors. The floor plan feature is then sent to the transformer backbone as a conditioning vector.
Transformer decoder. The backbone of our denoising network is a time-variant transformer decoder adapted from VQ-Diffusion [13]. Within each transformer block, the time step is injected through the Adaptive Layer Normalization (AdaLN) [3] operator. The conditioning inputs are sent through a multi-head cross-attention layer. Via this transformer decoder, the discrete semantic and continuous geometric features are jointly processed through attention to predict the next diffusion update.
Feature extractor. Finally, we feed the output of the transformer decoder to two MLPs to predict a categorical distribution for the semantic labels and Gaussian means for the geometric features respectively.
@@ 4.3 Partial Condition on Objects
3D scene synthesis is sometimes constrained by known object information, as for example in scene completion a given set of furniture pieces. DiffuScene [46] considers these as additional condition inputs to the denoising network, which requires at least a task-specific feature encoder for the conditioning inputs and hence also re-training of the entire model. We instead propose to address this issue through a corruption-and-masking strategy which does not require any additional training. Since the core idea of diffusion models is to recover real data by learning the inverse of a known corruption process, we can force the constrained sections of the latent variables to follow exactly the inverse of the controlled corruption steps in the denoising process. For example, for scene completion: given $M$ objects, we first compute a sequence of latent variables $\{(z_{t}^{i},{\bm{x}}_{t}^{i})\}_{1\leq i\leq M}$ through the corruption process. This allows us to mask out the first $M$ objects using pre-computed latent variables at time $t$ in the reverse process. These latent variables eventually converge to the known input $\{(z_{0}^{i},{\bm{x}}_{0}^{i})\}_{1\leq i\leq M}$ without additional training. This strategy can generalize to other applications, such as furniture arrangement and label conditioned scene synthesis, by targeting different sections of the latent variables.
@@ 5 Experimental Evaluation
In this section, we present evaluation results comparing MiDiffusion against existing state-of-the-art methods and provide ablation studies, demonstrating the usefulness of our contributions. We also show that, without any re-training, our models can be applied to other tasks such as scene completion and furniture arrangement. Further results and implementation details can be found in the appendix.
@@
Dataset. We use the common 3D-FRONT [10] benchmark for evaluation. 3D-FRONT is a synthetic dataset of 18,797 rooms furnished by 7,302 textured objects from 3D-FUTURE [11]. We evaluate on three room types: 4041 bedrooms, 900 dining rooms, and 813 living rooms. We use the same data processing and train/test data split as [32]. We use rotation augmentations in $90^{\circ}$ increments during training as in  [46] to be consistent with the baselines.
@@
Baselines. We compare our proposed approaches against two state-of-the-art 3D scene synthesis methods: (1) ATISS [32], an autoregressive model with a transformer encoder backbone that generates sequential predictions of objects; (2) DiffuScene [46], a diffusion model with a conditional 1D U-Net [38] denoising backbone based on DDPM. For each baseline model we use their open-source code and default hyper parameters, which includes a pretrained ResNet-18 [14] model as the floor plan feature extractor. For completeness, we additionally compare MiDiffusion with DiffuScene in its original application, i.e. unconditioned scene synthesis, in Appendix E.
@@
Implementation. We train all our models using the Adam optimizer [24] with a learning rate of $l_{r}=2e^{-4}$ and a $0.5$ decay rate every $10k$ for $50k$ epochs on the bedroom dataset, and every $15k$ for $100k$ epochs on the dining room and living room datasets. We set $\lambda=0.05$ for the auxiliary loss term $L_{aux}^{D3PM}$. More details about implementation and network parameters are presented in Appendix C.
@@
Evaluation Metrics. We randomly select 1000 floor plans from the test set as input and evaluate all models using common metrics from prior works [49, 32, 46]: KL divergence of object label distributions (KL×$\times 0.01$), Fréchet inception distance (FID), Kernel inception distance (KID×$\times 0.001$), and classification accuracy (CA%) following the implementation by ATISS. While the KL divergence evaluates the predicted object label distributions, the other metrics are used to compare the layout images with the ground-truth test data. We render each scene to a 256 x 256 image using CAD models from 3D-FUTURE through top-down orthographic projection. Each image captures a 6.2m square for bedrooms, or a 12.2m square living or dining room. In line with DiffuScene, we color each object according to its semantic class. Appendix D includes example layout images. We report CA % over 10 runs with random sampling of predicted layouts for training and testing. Note that realistic layouts should have a low FID/KID, a low KL and a CAthat is close to 50%. We also compute additional metrics, averaged over all scenes to further evaluate object placement, including the number of generated objects (Obj), percentage of out-of-boundary objects (OOB %), and bounding box intersection over union (IoU %). We dilate the room boundary by 0.1m for OOB % computation to account for slight under estimation of floor boundary in the raw data.
@@ 5.1 Floor Plan Conditioned 3D Scene Synthesis
In this section, we first compare MiDiffusion against the state-of-the-art in floor-conditioned 3D scene synthesis experiments without existing objects, before showing several ablations that highlight the impact of our proposed mixed diffusion paradigm.
@@
Comparison with state-of-the-art. In Fig. 2, we present qualitative results rendered with PyVista. Note that the object meshes are retrieved from the 3D-FUTURE [11] dataset by finding the closest match in size given the predicted semantic category. We replace the original textures in the CAD models by a label-specific color for clarity. We include examples of orthographic projection images which are used for quantitative evaluations in the appendix. In general, our approach is able to generate realistic object arrangements whilst respecting the boundary constraints. ATISS generates slightly less precise object placements compared to the two diffusion based approaches, e.g. the table-chairs furniture set in the last example of Fig. 2 is not exactly symmetric. Both DiffuScene and ATISS show a higher tendency to place objects outside the floor boundary than ours. This problem is more prominent for DiffuScene, probably because the conditional U-Net architecture in the denoising network was not intended for learning boundary constraints.
We show quantitative results in Tables 1 and 2. Overall, our approach achieves substantially better performance in generating realistic layouts, especially in the dining room and living room datasets. There are large margins between our models and the baseline models for FID, KID×$\times 0.001$ and CA %, when comparing images of synthesized layouts. As for object label distributions, there are no large differences between the individual methods with respect to KL×$\times 0.01$. In Table 2, there is a trade-off between keeping objects inside the boundary (low OOB%) and maintaining minimal overlap (low IoU%). Although DiffuScene achieves the lowest IoU%, their models tend to place more objects outside the boundary than other methods, which explains their sub-optimal performance in Table 1. Overall, our approach remains superior in terms of object placement, ensuring consistency between objects and compatibility with floor boundaries.
@@
ATISS [32] 65.76 1.21 54.44 $\pm$ 3.69 0.94 41.06 10.54 64.06 $\pm$ 5.63 1.89 38.03 9.20 63.98 $\pm$ 4.60 1.55
DiffuScene [46] 66.45 1.05 58.70 $\pm$ 4.93 3.41 45.88 9.30 65.34 $\pm$ 6.55 1.55 47.91 11.75 73.96 $\pm$ 4.93 1.95
MiDiffusion (Ours) 63.71 0.24 53.65 $\pm$ 2.46 1.05 30.63 2.74 52.10 $\pm$ 1.72 1.90 28.60 1.64 54.04 $\pm$ 3.33 1.54
@@
ATISS [32] 5.52 16.26 1.00 11.43 17.43 1.86 13.00 17.39 1.49
DiffuScene [46] 5.05 8.06 0.41 10.90 28.16 0.56 11.79 33.75 0.43
MiDiffusion (Ours) 5.22 3.91 0.61 10.92 5.77 0.91 11.91 7.93 0.98
@@
Ablation study. We start by ablating our main contribution, i.e. joint discrete-continuous diffusion. To this end, we first compare MiDiffusion (Mixed + PointNet) to a naive 2-stage approach (2-stage + PointNet). In the first stage we predict a set of object labels through discrete diffusion, which we then feed to the second stage to estimate their geometric attributes conditioned on the obtained semantic labels. For a fair comparison, we employ the same denoising architecture as in Fig. 1 for both stages and only remove the unused feature encoder or extractor from each stage. The networks are trained independently. Overall, this 2-stage approach is less efficient and less mathematically principled, and therefore obtains significantly worse results across all metrics compared to MiDiffusion in Table 3.
Next, we provide three variants of the joint prediction approach using the same transformer decoder backbone to show the effects of mixed diffusion against DDPM, and PointNet floor plan feature extractor against ResNet. Note that the first variant, DDPM+ResNet, only differs from DiffuScene in its transformer backbone and it already surpasses DiffuScene in Table 1, suggesting that the transformer architecture is better at learning boundary than its U-Net counterpart. When using the same floor plan extractor, the mixed diffusion variant always achieves better performance than the DDPM variant, stressing the importance of mixed diffusion. Further, while Mixed+ResNet uses the same floor feature extractor as the baselines in Table 1, it results in better performance. The more lightweight PointNet floor feature extractor brings a significant performance boost in the dining and living room datasets compared to ResNet, but results in slightly worse performance in the bedroom dataset. Since there are much less training scenes for dining and living rooms, and these scenes are larger in physical size, it is likely that PointNet works better under complex and limited training data while suffering from sampling noise during conversion of the floor plan image to boundary points and normals. Overall, we achieve the best results with our complete formulation (Mixed+PointNet) in which the mixed diffusion approach avoids lifting the discrete inputs to a continuous space as required by the DDPM formulation. We include additional evaluation results in Appendix E
Method Bedroom Dining room Living room
2-stage + PointNet 65.54 1.53 57.83 $\pm$ 4.89 3.50 38.80 7.02 61.53 $\pm$ 7.85 5.08 35.35 6.30 61.97 $\pm$ 6.55 4.73
Mixed+ResNet 63.88 0.20 54.41 $\pm$ 4.03 2.66 31.29 3.81 53.41 $\pm$ 3.48 2.04 30.15 2.81 57.21 $\pm$ 3.97 1.74
Mixed+PointNet 63.71 0.24 53.65 $\pm$ 2.46 1.05 30.63 2.74 52.10 $\pm$ 1.72 1.90 28.60 1.64 54.04 $\pm$ 3.33 1.54
Table 3: Ablation study for floor-conditioned 3D scene synthesis.
Diversity. We study the diversity of the predicted layouts by measuring the average standard deviation (std) of object centroid positions and bounding box sizes. Since vertical positions are highly correlated with semantic labels, which are already evaluated through KL-divergence, we compute the average std over the two planar axes only. The results are shown in Table 4, where the “IB” suffix means results are computed over “in-boundary” objects only to remove bad object placement. All three approaches are quite similar in terms of diversity. Overall, our approach generates the most diverse object placements that are within the floor boundary (i.e. Position-IB), but is slightly worse than DiffuScene in size variety.
@@
ATISS [32] 1.074 1.035 0.709 0.663 1.586 1.504 0.434 0.385 1.730 1.663 0.442 0.409
DiffuScene [46] 1.073 1.059 0.718 0.691 1.596 1.427 0.454 0.423 1.752 1.575 0.482 0.439
MiDiffusion (Ours) 1.073 1.067 0.698 0.680 1.568 1.558 0.414 0.399 1.668 1.667 0.439 0.428
Ground truth 1.091 1.082 0.722 0.690 1.623 1.618 0.450 0.447 1.735 1.729 0.463 0.460
@@
1 obj ATISS [32] 60.59 0.19 1.04 13.52 36.63 7.77 1.79 15.30 34.34 7.15 1.46 16.19
DiffuScene-SC [46] 83.90 19.17 18.34 3.11 34.12 4.82 1.59 12.13 36.32 6.64 1.30 21.74
MiDiffusion (ours) 60.08 -0.03 0.91 3.32 28.56 1.88 1.58 5.29 27.86 1.98 1.49 7.81
3 obj ATISS [32] 51.83 -0.57 0.66 8.89 33.53 5.39 1.60 12.26 31.78 5.39 1.43 13.03
DiffuScene-SC [46] 42.77 -1.08 0.80 5.13 30.58 3.67 0.84 10.04 32.07 4.58 0.57 16.13
MiDiffusion (ours) 52.19 -0.32 1.00 4.05 27.60 1.85 1.72 3.97 27.45 1.91 0.98 6.76